# DOCE: Finding the Sweet Spot for Execution-Based Code Generation

### v1

@@ Abstract
Recently, a diverse set of decoding and reranking procedures have been shown effective for LLM-based code generation. However, a comprehensive framework that links and experimentally compares these methods is missing. We address this by proposing Decoding Objectives for Code Execution, a comprehensive framework that includes candidate generation, $n$-best reranking, minimum Bayes risk (MBR) decoding, and self-debugging as the core components. We then study the contributions of these components through execution-based evaluation metrics. Our findings highlight the importance of execution-based methods and the difference gap between execution-based and execution-free methods. Furthermore, we assess the impact of filtering based on trial unit tests, a simple and effective strategy that has been often overlooked in prior works. We also propose self-debugging on multiple candidates, obtaining state-of-the-art performance on reranking for code generation. We expect our framework to provide a solid guideline for future research on code generation. [^1] Our code is available at https://github.com/deep-spin/doce.
@@ 1 Introduction
Despite the impressive capabilities of large language models (LLMs) in code generation (Chen et al., 2021; Li et al., 2022, 2023; Rozière et al., 2024; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024), demonstrated by their strong performance on popular execution-based metrics (Hendrycks et al., 2021; Chen et al., 2021; Austin et al., 2021; Lai et al., 2022), the predominant methods for generating code answers still rely on deterministic decoding techniques, such as greedy decoding and beam search, which can lead to suboptimal output (Shi et al., 2022).
Figure 1: Under the best settings, our proposed framework shows considerable improvement in execution accuracy over greedy decoding. We show results using CodeLlama-{7,13}B-Instruct (Rozière et al., 2024) and DeepSeekCoder-6.7B-Instruct (Guo et al., 2024a).
@@
An alternative approach involves generating multiple candidates through sampling a generative model and selecting one using $n$-best reranking (Ng et al., 2019; Bhattacharyya et al., 2021; Cobbe et al., 2021; Shen et al., 2021) or Minimum Bayes Risk (MBR) decoding (Eikema and Aziz, 2020, 2022). Metrics for reranking have been extensively explored in domains like machine translation (Eikema and Aziz, 2020; Freitag et al., 2022; Farinhas et al., 2023), where it has been shown to outperform greedy decoding and beam search. Similar methods in code generation include reranking with likelihoods of generated programs and instructions (Zhang et al., 2022) and MBR decoding with execution outputs (Shi et al., 2022). However, a unified framework providing an apples-to-apples comparison between these methods is currently missing.
@@
Figure 2: The Decoding Objectives for Code Execution (DOCE) Framework. Firstly, multiple candidates are generated through sampling. Each candidate then is assigned a score using an $n$-best reranker or MBR, before the candidate with the highest score is returned. Optionally, self-debugging can be applied to multiple candidates before scoring as we propose, or the highest score candidate as proposed by Chen et al. (2024).
Bridging this gap is especially significant for current research. First of all, with lately released open-source LLMs like CodeLlama (Rozière et al., 2024), there is demand for understanding the effects of reranking methods. For example, filtering based on trial unit tests (Li et al., 2022) was initially proposed when an LLM can generate few sensible code candidates, while lately released LLMs like CodeLlama (Rozière et al., 2024) can generate sensible programs easily.Additionally, evaluation with substantially more high-quality unit tests (Liu et al., 2023) is more rigorous and can potentially boost execution-based MBR decoding (Li et al., 2022; Shi et al., 2022). Last but not least, more metrics continue to emerge (Zhang et al., 2022; Zhou et al., 2023; Dong et al., 2023), providing an opportunity to comprehensively understand the execution component of reranking and self-debugging compared with execution-free approaches.
@@
Therefore, we propose a unified framework for execution-based code generation to improve decoding performance on execution-based metrics (see Figure 1). Our contributions are listed below:
[…]
@@
  • We analyze the impact of the number of generated candidates and sampling temperature. Specifically, sampling with previously unseen high temperatures allows high oracle reranking performance and execution-based MBR performance with filtering based on trial unit tests.
@@
  • With analysis of different approaches, we emphasize the importance of execution-based approaches in the context of candidate generation, reranking, and self-debugging. Specifically, our analysis reveals the importance of filtering based on trial unit tests, a commonly used technique whose effect has not been reported in previous works.
  • By applying DOCE to state-of-the-art language models, we find that execution-based reranking yields close-to-oracle performance. We show that self-debugging, a technique applied by Chen et al. (2024) as a final step, is more effective and robust when applied to multiple generated candidates before the final choice is made.
@@ 2.1 Candidate generation
An LLM-based code generation model defines a probability distribution $p_{\theta_{\mathrm{LLM}}}(y|x)$ over the space of possible programs, conditioned on a description of the task $x$, where $\theta_{\mathrm{LLM}}$ refers to learned parameters of the LLM. An $x$ often contains a natural language description; some unit test cases, namely trial unit tests, are used to regularize and generate the corresponding code, and the function heads optionally.
Generated code is usually predicted using greedy decoding or $p$-nucleus sampling (Holtzman et al., 2020). While greedy decoding generates a single high-probability candidate, sampling allows obtaining a set of diverse generated candidates $\mathcal{Y}$. Also note that in practice, nucleus sampling is often accompanied by temperature sampling (Dabre and Fujita, 2021). Previous studies (Shi et al., 2022; Chen et al., 2023) also found that sampling-based methods yield higher oracle performance (the oracle on models’ performance, choosing the optimal candidate) denoted by pass@k where k is the number of candidates generated). However, it is still unclear how to best pick a single candidate from the set of candidates.
@@ 2.2 Reranking
We now assume access to $\mathcal{Y}$ containing $N$ generated candidates for a given $x$, as mentioned in Section 2.1. To provide a final output $\hat{y}$, we explore how to rank and pick candidates for code generation effectively. Note that with outputs that have the same score, we choose the output with the smallest index.
@@
In its simplest form, $n$-best reranking uses a single feature $f$ and ranks candidates according to it, with the chosen candidate then being the one that maximizes this feature.
@@
Other than execution-based features, we can use a single model to predict the quality score of generated code with a trained model $\theta$. Specifically, previous efforts include training models using execution success/failure as signals (Inala et al., 2022; Ni et al., 2023), given the input $x$ and generated candidate $y^{\prime}$. During inference, a score $f_{\theta}(x,y^{\prime})$ is given for every generated candidate. Notice that for fair comparison of execution-based and execution-less methods, we do not consider models that require information from execution during inference time (Ni et al., 2023).
@@ 2.2.2 MBR Decoding
While $n$-best list reranking can rank candidates based on reference-free features, it is also possible to rank candidates based on reference-based metrics using MBR decoding. Concretely, if we consider a utility function $U(y^{\star},y)$ that measures the similarity between a candidate $y$ and some reference (correct) code $y^{\star}$, MBR selects the candidate in $\mathcal{Y}$ that maximizes the expected utility, when considering all candidates as possible references:
@@
While MBR is most frequently used to optimize translation quality metrics like BLEU (Papineni et al., 2002) or COMET (Rei et al., 2020), we are interested in metrics that specifically capture the quality of code. Shi et al. (2022) showed that translation-quality metrics do not outperform greedy decoding when evaluated with execution, for which we do not consider using translation-quality metrics like BLEU.
@@
In practice, this utility function reduces the MBR decision to a majority vote for the candidate that produces outputs that agree the most with other candidates’ outputs (Wang et al., 2023).
Note that in Shi et al. (2022) execution-based MBR both on the cases of using only inputs from trial unit tests only and from all unit tests used for evaluation. Since we decide to use these unit tests as filtering, we use all unit tests given but only their inputs. Our practice aligns with most previous common practices (Shi et al., 2022; Ni et al., 2023; Chen et al., 2024).
@@
We can also use external models for MBR. Existing external models include CodeBertScore (Zhou et al., 2023), CodeScore (Dong et al., 2023) to calculate $U(y,y^{\prime})$. Specifically, with a trained model $\theta$, we calculate $U_{\theta}(x,y,y^{\prime})$ by inputting input $x$ and two generated candidates $y$ and $y^{\prime}$.
@@ 2.3 Self-Debugging
The performance of reranking is oracleed by the maximum quality of generated candidates. An approach to further improve the quality of generated code is to debug it using the same model used for generation (Chen et al., 2024). Given input $x$, generation $y^{\prime}$, and optionally information of execution on trial unit tests $T_{\text{trial}}$, we obtain a new generation $y^{\prime}_{d}$ using the code LLM parameterized by $\theta_{\mathrm{LLM}}$:
$y^{\prime}_{d}=\mathrm{debug}_{\theta_{\mathrm{LLM}}}(x,y^{\prime},T_{\text{trial}}),$ (5)
where $y^{\prime}_{d}$ represents a corrected version $y$. In practice, this self-debugging can be achieved by prompting the model to correct errors in the presented code, potentially based on the results from execution. We use greedy decoding to get the debugged code, same as Chen et al. (2024).
@@ 2.4 DOCE Framework
We can also perform self-debugging on all generated candidates before reranking, namely SD-Multi. We presume this outperforms self-debugging on one candidate since self-debugging aims to regenerate candidates that pass the trial unit tests, and more candidates passing trial unit tests leads to more candidates for effective MBR decoding with possibly better reranking results. Therefore, the final code answer is $\operatorname*{argmax}_{y^{\prime}_{d}\in\mathcal{Y}_{d}}\frac{f(y^{\prime}_{d})}{|\mathcal{Y}_{d}|}\sum_{y_{d}\in\mathcal{Y}_{d}}U(y_{d},y^{\prime}_{d})$.
@@ 3.1 Datasets
We conduct experiments using two widely recognized execution-based datasets: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). We employ the sanitized, higher-quality subset of the MBPP dataset, namely MBPP-S. Note that HumanEval is a test-only dataset and MBPP-S are normally treated as test-only. We utilize EvalPlus (Liu et al., 2023) which includes more than 35 times unit tests than the original benchmark for both HumanEval and MBPP-S, ensuring a more rigorous assessment of function correctness.
@@ 3.2 Candidate Generation
We generate candidates with the CodeLlama-{7B,13B}-Instruct (Rozière et al., 2024) models and DeepSeekCoder-6.7B-Instruct (Guo et al., 2024a) model because instruction fine-tuning allows them to generate output in the markdown format with code blocks bracketed by ```python and ```, allowing self-debugged code to be processed easily. We generate {1,5,10,20,25,50} [^3] Note that we use candidate obtained with greedy decoding when with one generated candidate. candidates for each task. We first explore sampling temperature by varying it between 0.2 and 2.0 with $p$-nucleus sampling $p$ 0.95, before deciding what sampling temperature we use for subsequent experiments. With MBPP-S, we vary temperature between 0.2 and 1.8. When generating multiple candidates, we utilize the vLLM framework (Kwon et al., 2023) that enables fast inference. To answer research questions related to candidate generation and reranking, we present the results from CodeLlama-7B-Instruct.
@@ 3.3 Reranking
For $n$-best reranking, we analyze three metrics types: likelihood-based, execution-based, and neural. For likelihood-based metrics, we use the log-likelihood of the generated generations given the input $x$ as well as CoderReviewer (Zhang et al., 2022) using $\log(p(x|y))+\log(p(y|x))$. [^4] In practice, we use the likelihood normalized by sequence length, as it mitigates bias towards shorter sequences. For execution-based metrics, we consider passing the trial unit test case as mentioned in Equation 2. For neural metrics, we use CodeScore (Dong et al., 2023) as it is so far the only publicly available neural metric that does not rely on execution during inference. [^5] We only test CodeScore on HumanEval since it was also trained on the original MBPP dataset. We do not consider LEVER (Ni et al., 2023) as it requires execution during inference time and requires the reranker to be trained on the target task.
@@
For MBR, we consider execution-based metrics and neural metrics. Specifically, among execution-free metrics, we include CodeBertScore (Zhou et al., 2023), which is not trained with execution-based signals, and CodeScore (Dong et al., 2023), which is trained with execution-based signals. For execution-based metrics, we consider MBR-Exec (Shi et al., 2022) mentioned in Equation 4. Note that we assume access to the inputs of unit test cases used for evaluation, which aligns with one setting of MBR-Exec (Shi et al., 2022) and the setting of LEVER (Ni et al., 2023) and Self-Debug (Chen et al., 2024). We further study the effect of the number of test cases using EvalPlus (Liu et al., 2023), which contains over 100 high-quality unit test cases. Our setting is different from the setting using generated unit test cases (Chen et al., 2023; To et al., 2024; Huang et al., 2024) where normally neither trial unit test cases nor inputs of test cases for evaluation are available for either prompting or reranking. We leave the unification of these two settings to future work.
3.4 Self-Debugging
@@
We consider the simple setting of self-debugging proposed by Chen et al. (2024). We consider the setting of using only unit test (UT) feedback, i.e., the feedback obtained from the compiler when a generated candidate is tested on trial unit tests. We only consider this simple setting because 1) our main focus is to test generally the impact of self-debugging on improving the oracle’s performance, 2) it does not require steps of generation using the LLM other than self-debugging, thus requires less computation for inference and 3) it is the feedback that gives the largest gain to execution accuracy post-debugging according to Chen et al. (2024). We perform 3 rounds of self-debugging for candidates generated with sampling because Chen et al. (2024) demonstrated that most debugging can be finished in 3 rounds. In each round, the inputs to the LLM include few-shot examples, the refined candidate from the previous round, and the candidate’s execution feedback on trial unit tests.

### v4

@@ Abstract
A diverse set of decoding, reranking, and self-improvement procedures have been shown effective for code generation with LLMs. However, a unified framework that links and experimentally compares these methods is missing. We address this by proposing Decoding Objectives for Code Execution, a framework with candidate generation, $n$-best reranking, minimum Bayes risk (MBR) decoding, and self-debug as the core components. We then systematically study the contributions of these components through execution-based evaluation. Our findings suggest the importance of sampling with high temperature, reranking with execution-based methods using only a few high-quality unit test inputs, and robust self-debugging on multiple candidates. Specifically, we highlight filtering based on trial unit tests, a simple and effective strategy that has often been overlooked in prior works [^1] Our code is available at https://github.com/deep-spin/doce.
@@ 1 Introduction
Despite impressive performances of large language models (LLMs) in code generation (Chen et al., 2021; Li et al., 2022, 2023; Rozière et al., 2023; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024; DeepSeek-AI et al., 2024) on execution-based metrics (Hendrycks et al., 2021; Chen et al., 2021; Austin et al., 2021; Lai et al., 2023; Jain et al., 2024; Zhuo et al., 2024) based on exploration of training approaches, predominant methods during inference still heavily rely on techniques such as greedy decoding, beam search, and $p$-nucleus sampling (Holtzman et al., 2020) followed directly by evaluation on execution accuracy.
@@
An alternative approach involves sampling multiple candidates with an LLM and selecting one using $n$-best reranking (Ng et al., 2019; Bhattacharyya et al., 2021; Cobbe et al., 2021; Shen et al., 2021) or Minimum Bayes Risk (MBR) decoding (Eikema and Aziz, 2020, 2022), with potential improvements on candidates using the same LLM (Madaan et al., 2023). These methods have been systematically studied in domains like machine translation (Eikema and Aziz, 2020; Freitag et al., 2022; Fernandes et al., 2022; Farinhas et al., 2023). However, despite the proposal of reranking methods, based on likelihoods (Zhang et al., 2023), execution agreements (Shi et al., 2022) or predicting execution accuracy (Inala et al., 2022; Ni et al., 2023), and improvement methods simulating the debugging process (Olausson et al., 2024; Chen et al., 2024), a systematic study that helps understand the effects and provide an apples-to-apples comparison of these methods is still missing.
@@
Figure 1: The Decoding Objectives for Code Execution (DOCE) Framework. Firstly, multiple candidates are generated through sampling. Each candidate then is assigned a score using an $n$-best reranker or MBR, before the candidate with the highest score is returned. Self-Debug can be applied to multiple candidates before scoring as we propose, or the highest score candidate as proposed by Chen et al. (2024).
Bridging this gap is crucial to current research. First of all, lately released open-source code LLMs (Rozière et al., 2023; Luo et al., 2024; Guo et al., 2024a; DeepSeek-AI et al., 2024) excelling on a range of tasks lead to a demand for a renewed understanding of the effects of sampling and reranking methods. Example methods include choosing a relatively low sampling temperature (Li et al., 2022; Shi et al., 2022; Huang et al., 2024; Chen et al., 2023), which is originally suggested with LLMs having lower performances on these tasks (Fried et al., 2023; Nijkamp et al., 2023), and filtering on trial unit tests (Li et al., 2022), a method originally proposed to retain sensible programs when a code LLM can hardly generate. Additionally, strong open-source code LLMs also allow a systematic study of sampling, reranking, and candidate improvement with a low budget compared to models requesting API calls (OpenAI, 2023). Last but not least, lately proposed neural metrics for code (Zhang et al., 2023; Zhou et al., 2023; Dong et al., 2023) provide an opportunity to systematically understand the execution component of reranking compared with execution-free approaches.
@@
Therefore, we propose a unified framework for execution-based code generation to improve decoding performance on execution-based metrics. Our contributions are listed below:
  • We propose the Decoding Objectives for Code Execution (DOCE) framework for code generation (see Figure 1), including candidate generation, reranking, and self-debug.
@@
  • We provide analysis on the number of generated candidates and sampling temperature. Specifically, sampling with previously unseen high temperatures allows high oracle and reranking performance.
@@
  • With experiments on reranking, we emphasize the importance of execution-based approaches. Specifically, our analysis reveals the importance of filtering based on trial unit tests, a commonly used technique whose effect has been rarely reported in previous works.
  • Execution-based methods lead to close-to-Oracle performance, with further improvements using Self-Debug. Effective and robust Self-Debug only happens when applied to multiple candidates before choosing the final candidate.
@@ 2.1 Candidate generation
An LLM-based code generation model defines a probability distribution $p_{\theta_{\mathrm{LLM}}}(y|x)$ over the space of possible programs, conditioned on a description of the task $x$, where $\theta_{\mathrm{LLM}}$ refers to learned parameters of the LLM. An $x$ often contains a natural language description; some unit test cases, namely trial unit tests, are used to regularize and generate the corresponding code.
Generated code is usually predicted using greedy decoding or $p$-nucleus sampling (Holtzman et al., 2020). The latter approach allows obtaining diverse generated candidates $\mathcal{Y}$ compared to greedy decoding and is often accompanied by temperature sampling (Dabre and Fujita, 2021). Previous studies (Shi et al., 2022; Chen et al., 2023) found that sampling leads to higher oracle performance, denoted by pass@k, assuming always choosing the best among k candidates. However, deciding which candidate to pick is still unclear.
@@ 2.2 Reranking
We now assume access to $\mathcal{Y}$ containing $N$ generated candidates for a given $x$, as mentioned in Section 2.1. To provide a final output $\hat{y}$, we explore how to rank and pick candidates for code generation effectively [^2] With outputs that have the same score, we choose the output with the smallest index..
@@
In its simplest form, $n$-best reranking uses a single feature $f$ and ranks candidates, with the chosen candidate maximizing this feature:
@@
Likelihood-based Feature.
We can use the likelihood as the feature $f(y^{\prime})=p_{\theta_{LLM}}(y^{\prime}|x)$. In practice of code generation, the likelihood is often normalized by sequence length $p_{\theta_{LLM}}(y^{\prime}|x)^{\frac{1}{len(y^{\prime})}}$. Alternatively, Zhang et al. (2023) proposed an average of likelihood combined with a “reviewer-likelihood”, normalized by sequence length as well:
$f(y^{\prime})=(p_{\theta_{LLM}}(y^{\prime}|x)p_{\theta_{LLM}}(x|I,y^{\prime}))^{\frac{1}{len(y^{\prime})}},$ (2)
where $I$ is the set of the few-shot demonstrations.
@@
We can also use a single model $\theta_{ext}$. Specifically, previous efforts include training models using execution success/failure as signals (Inala et al., 2022; Ni et al., 2023), given the input $x$ and generated candidate $y^{\prime}$. During inference, a score $f_{\theta_{ext}}(x,y^{\prime})$ is given for every generated candidate.
@@ 2.2.2 MBR Decoding
While $n$-best list reranking can rank candidates based on reference-free features, it is also possible to rank candidates with reference-based metrics using MBR decoding. With a utility function $U(y^{\star},y)$ measuring the similarity between a candidate $y$ and some reference (correct) code $y^{\star}$ MBR selects the candidate in $\mathcal{Y}$ that maximizes the expected utility considering all candidates as possible references:
@@
In practice, this utility function reduces the MBR decision to a majority vote for the candidate that produces outputs that agree the most with other candidates’ outputs (Wang et al., 2023a).
Note that in Shi et al. (2022) execution-based MBR both on the cases of using only inputs from trial unit tests only and from all unit tests used for evaluation. Since we decide to use these unit tests as filtering, we use all unit tests given but only their inputs, which aligns with previous practices (Shi et al., 2022; Ni et al., 2023; Chen et al., 2024).
@@
We can also use an external neural model (Zhou et al., 2023; Dong et al., 2023) used for evaluation $\theta_{ext}$ to compute $U_{\theta_{ext}}(x,y,y^{\prime})$, obtaining the utility by inputting input $x$ and two generated candidates $y$ and $y^{\prime}$.
@@ 2.3 Self-Debugging
Reranking is oracleed by the best possible candidate’s performance. An approach for further quality improvement of generated codes is to debug it with the same model used for generation (Chen et al., 2024). Given few-shot demonstrations $I$, an input $x$, a candidate $y^{\prime}$, and its execution feedback with the trial unit tests $exec(y^{\prime},T_{\text{trial}})$ , we obtain a corrected generation $y^{\prime}_{d}$ using $\theta_{\mathrm{LLM}}$:
$y^{\prime}_{d}=\mathrm{debug}_{\theta_{\mathrm{LLM}}}(I,x,y^{\prime},exec(y^{\prime},T_{\text{trial}})).$ (6)
@@ 2.4 DOCE Framework
We can also perform self-debugging on all generated candidates before reranking, namely SD-Multi. We presume this outperforms self-debugging on one candidate with more candidates passing trial unit tests, leading to better candidates to be reranked and possibly better reranking results. The final code answer is $\operatorname*{argmax}_{y^{\prime}_{d}\in\mathcal{Y}_{d}}\frac{f(y^{\prime}_{d})}{|\mathcal{Y}_{d}|}\sum_{y_{d}\in\mathcal{Y}_{d}}U(y_{d},y^{\prime}_{d})$.
@@ 3.1 Datasets
We first conduct experiments using two widely recognized execution-based datasets: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). We employ the sanitized, higher-quality subset of the MBPP dataset, namely MBPP-S. We utilize EvalPlus (Liu et al., 2023) that includes more than 35 times unit tests than the original benchmark, ensuring rigorous execution-based evaluation. We then extend experiments to competitive programming using LiveCodeBench Jain et al. (2024), a dataset containing up-to-date competitive programming problems to avoid data contamination. See Appendix A for further details.
3.2 Code LLMs
We generate candidates with the CodeLlama-{7B,13B}-Instruct (Rozière et al., 2023) and DeepSeekCoder-{6.7B,V2-Lite}-Instruct (Guo et al., 2024a; DeepSeek-AI et al., 2024). We use instruction-tuned models that enable simple processing of generated codes for both candidate generation and Self-Debug.
@@ 3.3 Candidate Generation
We generate 5 to 50 candidates for each task. We first explore sampling temperature by varying it between 0.2 and 2.0 with $p$-nucleus sampling $p$ 0.95, before deciding what sampling temperature we use for subsequent experiments. With MBPP-S and LiveCodeBench, we vary temperature between 0.2 and 1.8. When generating multiple candidates, we use the open-source vLLM (Kwon et al., 2023) for fast inference. To answer research questions related to candidate generation and reranking, we present the results from CodeLlama-7B-Instruct.
@@ 3.4 Reranking
For $n$-best reranking, we study three metrics types: likelihood-based, execution-based, and neural. For likelihood-based metrics, we use the likelihood feature and CoderReviewer, both normalized by sequence length. For execution-based metrics, we consider filtering based on trial unit tests. For neural metrics, we use CodeScore (Dong et al., 2023) as it is so far the only publicly available general-purpose neural metric trained to predict execution accuracy. We do not consider task-specific and model-specific neural metrics (Inala et al., 2022; Ni et al., 2023).
@@
For MBR, we consider execution-based metrics and neural metrics. Specifically, among execution-free metrics, we include CodeBertScore (Zhou et al., 2023), which is not trained with execution signals, and CodeScore (Dong et al., 2023), which is trained with execution signals. For execution-based metrics, we consider MBR-Exec (Shi et al., 2022) mentioned in Equation 5. Note that we assume access to the inputs of unit tests used for evaluation, which aligns with one setting of MBR-Exec (Shi et al., 2022) and the setting of LEVER (Ni et al., 2023) and Self-Debug (Chen et al., 2024).
3.5 Self-Debug
@@
We consider the simple setting of Self-Debug proposed by Chen et al. (2024), using only unit test (UT) feedback, i.e., the feedback obtained from execution when a generated candidate is tested on trial unit tests. We only consider this simple setting because 1) our main focus is to test generally the impact of Self-Debug on improving the oracle’s performance, 2) it does not require steps of generation using the LLM other than Self-Debug, thus requires less computation for inference and 3) it is the feedback that gives the largest gain to execution accuracy post-debugging according to Chen et al. (2024). We perform 3 rounds of Self-Debug for candidates generated with sampling because Chen et al. (2024) demonstrated that most debugging can be finished in 3 rounds. In each round, the LLM debugs the candidates generated in the last round.
4 Results
In this section, we provide results and discuss the effects of sampling, reranking, and Self-Debug. Firstly, we analyze the impact of candidate generation on the model’s oracle performance (assuming we pick the best possible candidate) and the performance of execution-based reranking methods. We then compare different reranking methods and discuss the importance of execution for reranking. With the answers to these questions in mind, we study the effect of Self-Debug and discuss the best way to combine all these methods to improve reranking performance. We provide a subset of all our experimental results to answer these research questions. For results on more experiments, see Appendix B.
4.1 Candidate Generation
First, we study the impact of candidate generation varied by candidate numbers and sampling temperature using CodeLlama-7B-Instruct. We provide average execution accuracies, reranking results using MBR-Exec with and without filtering based on trial unit tests, and estimates of reranking oracle using Pass@k. We hypothesize that filtering gives reranking better alignment with the (potential) ground-truth function as it is a strong signal.